OpenStudy (anonymous):

How many radians on the unit circle would the minute hand travel from 0° if it were to move 5π inches?

OpenStudy (anonymous):

On the unit circle, wich is a circle of radius 1, it will move 2pi inches for 2pi radians. (because the entire circle's circumference is 2pi(radius) = 2pi(1) = 2pi). So you want 5pi inches. I think you can now complete the problem.

OpenStudy (anonymous):

I really dont understand this question though is how the entire circle would be 2pi but they want 5pi would 5pi be smaller than 2pi? are is it saying it went around the clock multiple times?

OpenStudy (anonymous):

you could use cross multiplication

OpenStudy (anonymous):

Keep in mind that if a circle has a radius of 1, then the length of the entire arc of a circle, which is its circumference, will be 2pi units. If the radius is in inches, then the entire circumference of the circle will be 2pi inches.

OpenStudy (anonymous):

Or you can recall the formula, S = r(theta), where S = length of the arc r = radius of the circle theta = # of radians in the central angle

terenzreignz (terenzreignz):

I guess the thing to remember is that for every \(2\pi\) radians the minute hand moves from zero degrees, it's basically back to zero.

terenzreignz (terenzreignz):

@helppls123 a drawing? 

terenzreignz (terenzreignz):

I marked out zero degrees, and suppose the minute hand points there initially. Now for every \(2\pi\) inches it travels, it returns to that zero-degree point, aye?

OpenStudy (skyz):

IMa cozy up into here

terenzreignz (terenzreignz):

@Skyz You -_-

OpenStudy (skyz):

well.. i feel loved

OpenStudy (anonymous):

okay so like i was saying it is like the hand went around the clock multiple times?

terenzreignz (terenzreignz):

@Skyz your fans at TC await you LOL @help123please. that's how I understand it :D How many \(2\pi\)'s go into \(5\pi\)?

OpenStudy (anonymous):

twice fully and half?

terenzreignz (terenzreignz):

How many times fully is all that matters, and that means twice fully. So after it runs \(2\pi\) twice, it's again at zero degrees, with \(\pi\) more radians to go.

OpenStudy (anonymous):

so would it be 4pi?

OpenStudy (anonymous):

or would it be 5pi?

terenzreignz (terenzreignz):

No, you misunderstand...after it travels 4pi inches, it's basically travelled 0 inches, since after 4pi, it's back where it started, leaving you with pi more inches to travel.

OpenStudy (mertsj):

If it moved 5 pi inches, it would go around the circle 2 1/2 times. Therefore it would move 5 pi radians.

OpenStudy (anonymous):

i think my brain is going to explode. okay so it wants me to move it from 0 to 5pi right?

terenzreignz (terenzreignz):

Whoops... I must have misunderstood the question -_- --awkward-- Thanks for sorting that out Mertsj ^_^

OpenStudy (anonymous):

Mertsj...agree.

OpenStudy (mertsj):

It says that a point moved 5 pi inches. If the radius is 1, the circumference is 2 pi so if it moved 5 pi that is 2 1/2 times around the circle. The number of radian in 1 revolution of any circle is 2 pi radians. So in 2 1/2 revolutions it is 5 pi radians.

terenzreignz (terenzreignz):

I concede. @helppls123 apologies, it was simpler than I anticipated D:

OpenStudy (anonymous):

So the answer is 2 1/2? o-o i'm sorry i'm so confused right now and don't worry terenz this stuff is all kinds of crazy lmao

OpenStudy (mertsj):

The answer is 5 pi

terenzreignz (terenzreignz):

No, when the unit circle is concerned, every one unit = one radian In this case, your unit is inches. Now, a unit circle is involved, the minute hand traveled 5pi inches, and since one inch is one radian, it would have traveled 5pi radians

OpenStudy (anonymous):

but that is littereally in the question o-o is it really that obvious?

terenzreignz (terenzreignz):

That's what threw me off too :>

OpenStudy (mertsj):